1ST OF an AP is unity and the ratio of the sum of the 1st five terms to the sum of the next five terms is 9:29.find the AP

Grade:12

1 Answers

Shashwat Tiwari
12 years ago

Sum of 1st 5 terms=S1=5/2x(2+4d)

sum of next five terms=S2=sum of first 10 terms-sum of first 5 terms..............

                                  =5(2+9d)-(5/2(2+4d)..........

 

So, S1/S2=9/29

you get d=4

So, the A.P. is, 1,5,9,13...................